OpenStudy (anonymous):

The mean of discrete random variance variable X is 19.59. What is the variance?

OpenStudy (anonymous):

@camerondoherty

OpenStudy (amistre64):

mean = np variance = np(1-p) but there are way to meany versions of np for the factors of 19.59 to determine that with from what i see ...

OpenStudy (kirbykirby):

Do you know what kind of random variable it is?

OpenStudy (amistre64):

lol, its just some random variable you find lying in the street :)

OpenStudy (anonymous):

(x,p(x)) = (12,0.07) (15,0.21) (17,0.17) (20,0.25) (22,0.05) (24,0.04) (25,0.13) (30,0.08)

OpenStudy (kirbykirby):

Well what is it's distribution. We can't assume the mean is np unless it's binomial

OpenStudy (anonymous):

discrete random variable

OpenStudy (amistre64):

hmm, i do like making assumptions, but ill concede this point

OpenStudy (kirbykirby):

\[ E(X^2)=\sum_{x}x^2p(x)\] and \[Var(X)=E(X^2)-[E(X)]^2 \]

OpenStudy (amistre64):

\[E(X)=\sum xp(x)\] \[VAR(X)=\sum [x-E(X)]^2p(x)\]

OpenStudy (amistre64):

yeah, something like that :)

OpenStudy (kirbykirby):

they are the same :)

OpenStudy (anonymous):

It says...The following table shows the probability distribution for a discrete variable. Then the chart i listed above, and then the question i asked

OpenStudy (amistre64):

we have presented you with a few ways to approach the solution .... any questions about these methods?

OpenStudy (anonymous):

I have no idea what that formula means

OpenStudy (anonymous):

or what to do with it?

OpenStudy (amistre64):

have you read your lesson materials?

OpenStudy (anonymous):

There is no lesson material

OpenStudy (anonymous):

it is an online credit recovery program for highschool and it just gives quizzes from algebra 2 that i am supposed to know how to do already

OpenStudy (amistre64):

can you use excel?

OpenStudy (amistre64):

the process is simple enough, its just time consuming with so many data points

OpenStudy (kirbykirby):

They give you a pair of (x, p(x)), so to find \[E(X^2)=\sum_{x}x^2p(x)\], you multiply the square of the x-value with p(x). Do this for every pair they gave you, then add them up. \[E(X^2)=12^2(0.07)+15^2(0.21)+...+30^2(0.08) \] The the variance is \(Var(X)=E(X^2)-[E(X)]^2\). Now \(E(X)\) is just the mean that they gave you above (19.59) So once you calculate \(E(X^2)\) from above, just do: \(Var(X)=E(X^2)-(19.59)^2\)